# Hadronic molecules with four charm or beauty quarks

### v1

@@ II.3 The $cc\bar{b}\bar{b}$ and $\bar{c}\bar{c}bb$ systems
$C^{t}_{PP}$ $=$ $\frac{1}{m^{2}_{J/\psi}}+\frac{1}{m^{2}_{\Upsilon}}\,,$ (29)
@@
The $VP$ interaction involves only one channel $B_{c}^{*+}B_{c}^{+}$, whose coefficient is
$C^{t}_{VP}$ $=$ $\frac{1}{m^{2}_{J/\psi}}+\frac{1}{m^{2}_{\Upsilon}}\,.$ (30)
@@
$C^{t}_{VV}$ $=$ $\frac{1}{m^{2}_{J/\psi}}+\frac{1}{m^{2}_{\Upsilon}}\,,$ (31)
$C^{u}_{VV}$ $=$ $\frac{1}{m^{2}_{J/\psi}}+\frac{1}{m^{2}_{\Upsilon}}\,,$
and the relevant contact term is
@@
The above coefficients do not support the existence of hadronic molecules in the $cc\bar{b}\bar{b}$ system.
II.4 The $bb\bar{c}\bar{b}$ and $\bar{b}\bar{b}cb$ systems
The results for the $bb\bar{c}\bar{b}$ and $\bar{b}\bar{b}cb$ systems are the same, so we only need to study the $bb\bar{c}\bar{b}$ system. In this system, the $PP$ interaction involves only one channel $\eta_{b}B_{c}^{-}$, whose coefficients are
$C^{t}_{PP}$ $=$ $0\,,$ (36)
$C^{u}_{PP}$ $=$ $\lambda\frac{1}{m^{2}_{B_{c}^{*}}}\,.$
Hence, our result suggests that there do not exist hadronic molecules in the $PP$ sector.
The $VP$ interaction involves two coupled channels $\Upsilon B_{c}^{-}$ and $B_{c}^{*-}\eta_{b}$, whose coefficient is
$C_{VP}^{t}$ $=$ $\left(\begin{array}[]{c|cc}J=1&\Upsilon B_{c}^{-}&B_{c}^{*-}\eta_{b}\\ \hline\cr\Upsilon B_{c}^{-}&0&\lambda\frac{1}{m_{B_{c}^{*}}^{2}}\\ B_{c}^{*-}\eta_{b}&\lambda\frac{1}{m_{B_{c}^{*}}^{2}}&0\end{array}\right)\,.$ (40)
$|VP^{\prime+}\rangle$ $=$ $\frac{1}{\sqrt{2}}\left(|\Upsilon B_{c}^{-}\rangle+|B_{c}^{*-}\eta_{b}\rangle\right),$ (41)
$|VP^{\prime-}\rangle$ $=$ $\frac{1}{\sqrt{2}}\left(|\Upsilon B_{c}^{-}\rangle-|B_{c}^{*-}\eta_{b}\rangle\right),$ (42)
whose coefficient is
$C_{VP}^{\prime t}$ $=$ $\left(\begin{array}[]{c|cc}J=1&VP^{\prime+}&VP^{\prime-}\\ \hline\cr VP^{\prime+}&\lambda\frac{1}{m_{B_{c}^{*}}^{2}}&0\\ VP^{\prime-}&0&-\lambda\frac{1}{m_{B_{c}^{*}}^{2}}\end{array}\right)\,.$ (46)
Hence, the interaction due to the exchange of the $B_{c}^{*}$ meson in the $VP^{\prime-}$ channel turns out to be attractive, so there may exist a hadronic molecule of $J^{P}=1^{+}$ in the $VP$ sector.
The $VV$ interaction involves only one channel $\Upsilon B_{c}^{*-}$, whose coefficients are
$C^{t}_{VV}$ $=$ $0\,,$ (47)
$C^{u}_{VV}$ $=$ $\lambda\frac{1}{m^{2}_{B_{c}^{*}}}\,,$
and the relevant contact term is
$V^{co}_{\Upsilon B_{c}^{*-}\to\Upsilon B_{c}^{*-}}(s)$ $=$ $\left\{\begin{array}[]{cc}-2g^{2}&~{}\textrm{for $J=0$},\\ 3g^{2}&~{}\textrm{for $J=1$},\\ g^{2}&~{}\textrm{for $J=2$}.\end{array}\right.$ (51)
II.5 The $bb\bar{b}\bar{b}$ system
In the $bb\bar{b}\bar{b}$ system, the $PP$ interaction involves only one channel $\eta_{b}\eta_{b}$, the $VP$ interaction involves only one channel $\Upsilon\eta_{b}$, and the $VV$ interaction also involves only one channel $\Upsilon\Upsilon$.
Since the two vertices are both zero, $\mathcal{L}_{\eta_{b}\eta_{b}\Upsilon}=0$ and $\mathcal{L}_{\Upsilon\Upsilon\Upsilon}=0$, the interactions in this system all vanish within the extended local hidden gauge framework:
$C_{PP}^{t}$ $=$ $0\,,$
$C_{PP}^{u}$ $=$ $0\,,$
$C_{VP}^{t}$ $=$ $0\,,$ (52)
$C_{VV}^{t}$ $=$ $0\,,$
$C_{VV}^{u}$ $=$ $0\,,$
$C_{VV}^{co}$ $=$ $0\,.$
Therefore, our results do not support the existence of hadronic molecules in the $bb\bar{b}\bar{b}$ system.
@@ III Numerical Results
In the previous section we have studied the interactions of the $cc\bar{c}\bar{c}$, $cc\bar{c}\bar{b}/\bar{c}\bar{c}cb$, $cc\bar{b}\bar{b}/\bar{c}\bar{c}bb$, $bb\bar{c}\bar{b}/\bar{b}\bar{b}cb$, and $bb\bar{b}\bar{b}$ systems. In this section we numerically study their properties. As shown in Eq. (10), the loop function $G(s)$ is regularized using the cutoff method, with the cutoff momentum $\Lambda$ describing the dynamical scale to be integrated out. Its value is quite uncertain for the exchange of fully-heavy vector mesons, and we follow Ref. Liu:2023gla to choose a broad region, $\Lambda=400\sim 1400\textrm{ MeV}$, to perform numerical analyses.
Within the extended local hidden gauge framework, the resonances are dynamically generated as poles of the scattering amplitudes $T_{PP/VP/VV}(s)$. We find the existence of eight poles that may lead to some singular structures on the invariant mass spectrum:
  • There exists a pole of $J^{P}=1^{+}$ generated in the $VP$ sector of the $cc\bar{c}\bar{b}$ system.
  • There exists a pole of $J^{P}=1^{+}$ generated in the $VV$ sector of the $cc\bar{c}\bar{b}$ system.
  • There exists a pole of $J^{P}=1^{+}$ generated in the $VP$ sector of the $bb\bar{c}\bar{b}$ system.
  • There exists a pole of $J^{P}=1^{+}$ generated in the $VV$ sector of the $bb\bar{c}\bar{b}$ system.
We summarize their positions in Table 2 with respect to the cutoff momentum $\Lambda$. Besides, there exist four charge-conjugated poles in the $\bar{c}\bar{c}cb$ and $\bar{b}\bar{b}cb$ systems.
@@
We find two poles in the $cc\bar{c}\bar{b}$ system: one pole in the $VP$ sector and the other in the $VV$ sector. However, both of them correspond to the virtual states when setting the cutoff momentum $\Lambda=400\sim 1400\textrm{ MeV}$, so they can only result in some threshold effects. Actually, the pole in the $VP$ sector corresponds to the sub-threshold bound state only with Λ>1550 $\Lambda>1550\textrm{ MeV}$, and the pole in the $VV$ sector corresponds to the sub-threshold bound state only with Λ>2650 $\Lambda>2650\textrm{ MeV}$. Therefore, our results do not support the existence of deeply-bound hadronic molecules in the $cc\bar{c}\bar{b}$ system.
@@ IV Conclusion
In this paper we systematically study the fully-heavy meson-meson interactions with the quark constituents $cc\bar{c}\bar{c}$, $cc\bar{c}\bar{b}/\bar{c}\bar{c}cb$, $cc\bar{b}\bar{b}/\bar{c}\bar{c}bb$, $bb\bar{c}\bar{b}/\bar{b}\bar{b}cb$, $bb\bar{b}\bar{b}$ through the extended local hidden gauge formalism. After solving the coupled-channel Bethe-Salpeter equation, we search for poles on both the first (physical) and second Riemann sheets. The obtained results are summarized in Table 2 with respect to the cutoff momentum $\Lambda$.
We find two poles in the $bb\bar{c}\bar{b}$ system: the pole generated in the $VP$ sector corresponds to the sub-threshold bound state when setting the cutoff momentum Λ>850 $\Lambda>850\textrm{ MeV}$, while it causes the threshold effect with Λ<850 $\Lambda<850\textrm{ MeV}$; the pole generated in the $VV$ sector corresponds to the sub-threshold bound state with Λ>1100 $\Lambda>1100\textrm{ MeV}$. These two poles are potential fully-heavy hadronic molecules, and we propose to investigate them in the $\mu^{+}\mu^{-}B_{c}^{-}$ channel at LHC. We also find two poles in the $cc\bar{c}\bar{b}$ system: the pole generated in the $VP$ sector corresponds to the sub-threshold bound state only with Λ>1550 $\Lambda>1550\textrm{ MeV}$, and the pole generated in the $VV$ sector corresponds to the sub-threshold bound state only with Λ>2650 $\Lambda>2650\textrm{ MeV}$. Besides, we find four charge-conjugated poles in the $\bar{c}\bar{c}cb$ and $\bar{b}\bar{b}cb$ systems. However, our results do not support the existence of hadronic molecules in the $cc\bar{c}\bar{c}$, $cc\bar{b}\bar{b}/\bar{c}\bar{c}bb$, and $bb\bar{b}\bar{b}$ systems.

### v2

@@ II.3 The $cc\bar{b}\bar{b}$ and $\bar{c}\bar{c}bb$ systems
$C^{t}_{PP}$ $=$ $\frac{1}{m^{2}_{J/\psi}}+\frac{1}{m^{2}_{\Upsilon}}\,\quad and$ (31)
@@
The $VP$ interaction involves only one channel $B_{c}^{*+}B_{c}^{+}$, whose coefficient are
$C^{t}_{VP}$ $=$ $\frac{1}{m^{2}_{J/\psi}}+\frac{1}{m^{2}_{\Upsilon}}\,\quad and$ (32)
$C_{VP}^{u}$ $=$ $0\,.$
@@
$C^{t}_{VV}$ $=$ $\frac{1}{m^{2}_{J/\psi}}+\frac{1}{m^{2}_{\Upsilon}}\,\quad and$ (33)
$C^{u}_{VV}$ $=$ $\frac{1}{m^{2}_{J/\psi}}+\frac{1}{m^{2}_{\Upsilon}}\,.$
The relevant contact term is
@@
In sectors other than the $|(VV)_{cc\bar{b}\bar{b}};J^{P}=0^{+}\rangle$ sector, the coefficients are positive, corresponding to repulsive interactions. As for the $|(VV)_{cc\bar{b}\bar{b}};J^{P}=0^{+}\rangle$ sector, although the contact term provides an attractive potential of $-8g^{2}$, the repulsive potential generated by the exchange of vector mesons is approximately $+13g^{2}$ at threshold, resulting in an overall repulsive interaction. Therefore, in these sectors, the exchange of vector mesons cannot bind mesons together, hence the above coefficients do not support the existence of hadronic molecules in the $cc\bar{b}\bar{b}$ system.
@@ III Numerical Results
In the previous section we have studied the interactions of the $cc\bar{c}\bar{c}$, $cc\bar{c}\bar{b}/\bar{c}\bar{c}cb$, $cc\bar{b}\bar{b}/\bar{c}\bar{c}bb$, $bb\bar{c}\bar{b}/\bar{b}\bar{b}cb$, and $bb\bar{b}\bar{b}$ systems. In this section we numerically study their properties. As shown in Eq. (10), the loop function $G(s)$ is regularized using the cutoff method, with the cutoff momentum $\Lambda$ describing the dynamical scale to be integrated out. Its value is quite uncertain for the exchange of fully-heavy vector mesons, and we follow Ref. Liu:2023gla to choose a broad region, $\Lambda=400\sim 1400\textrm{ MeV}$, to perform numerical analyses, since the authors of Refs. Lu:2014ina ; Ozpineci:2013zas have already found that the requirement of heavy quark symmetry demands the use of the same cutoff momentum in the charm and bottom sectors. We note that the value of this important parameter is quite uncertain for the exchange of fully-heavy vector mesons, so the present study as well as Ref. Liu:2023gla serve as pioneer researches to investigate the fully-heavy hadronic molecules, but there do exist large theoretical uncertainties.
Within the extended local hidden gauge framework, the resonances are dynamically generated as poles of the scattering amplitudes $T_{PP/VP/VV}(s)$. We find the existence of eight poles that may lead to some singular structures on the invariant mass spectrum. There exist two poles in the $cc\bar{c}\bar{b}$ system and two poles in the $bb\bar{c}\bar{b}$ system, which we shall discuss in details later. We summarize their positions in Table 2 with respect to the cutoff momentum $\Lambda$. Besides, there exist four charge-conjugated poles in the $\bar{c}\bar{c}cb$ and $\bar{b}\bar{b}cb$ systems.
@@
We find two poles in the $cc\bar{c}\bar{b}$ system: one pole in the $VP$ sector and the other in the $VV$ sector. However, both of them correspond to virtual states when setting the cutoff momentum $\Lambda=400\sim 1400\textrm{ MeV}$, so they can only result in some threshold effects. The pole in the $VP$ sector corresponds to the sub-threshold bound state with Λ>1550 $\Lambda>1550\textrm{ MeV}$, and the pole in the $VV$ sector corresponds to the sub-threshold bound state with Λ>2650 $\Lambda>2650\textrm{ MeV}$. We generally consider the cutoff momentum $\Lambda$ to be consistent with the chiral unitary approach, which takes $\Lambda\approx 4\pi f_{\pi}\simeq 1200$ MeV. This value reflects certain non-perturbative properties of QCD. However, it is important to note that the parameter $\Lambda$ also functions as a free parameter, absorbing some implicitly considered interactions, and as a result, it may deviate from 1200 MeV in practical applications. Empirically, $\Lambda$ is usually taken to fall within the range of $400$ to $700$ MeV. Given the uncertainties inherent in our work, we believe it is appropriate to extend this range to $400$ to $1400$ MeV. Therefore, our results do not support the existence of deeply-bound hadronic molecules in the $cc\bar{c}\bar{b}$ system.
@@ IV Conclusion
In this paper we study the fully-heavy meson-meson interactions with the quark constituents $cc\bar{c}\bar{c}$, $cc\bar{c}\bar{b}/\bar{c}\bar{c}cb$, $cc\bar{b}\bar{b}/\bar{c}\bar{c}bb$, $bb\bar{c}\bar{b}/\bar{b}\bar{b}cb$, $bb\bar{b}\bar{b}$ through the extended local hidden gauge formalism. After solving the coupled-channel Bethe-Salpeter equation, we search for poles on both the first (physical) and second Riemann sheets. The obtained results are summarized in Table 2 with respect to the cutoff momentum $\Lambda$. We find two poles in the $bb\bar{c}\bar{b}$ system (and two charge-conjugated poles in the $\bar{b}\bar{b}cb$ system): the pole generated in the $VP$ sector corresponds to the sub-threshold bound state when setting the cutoff momentum Λ>850 $\Lambda>850\textrm{ MeV}$, and the pole generated in the $VV$ sector corresponds to the sub-threshold bound state with Λ>1100 $\Lambda>1100\textrm{ MeV}$. These two poles are potential fully-heavy hadronic molecules, and we propose to investigate them in the $\mu^{+}\mu^{-}B_{c}^{-}$ channel at LHC. However, our results do not support the existence of hadronic molecules in the $cc\bar{c}\bar{c}$, $cc\bar{c}\bar{b}/\bar{c}\bar{c}cb$, $cc\bar{b}\bar{b}/\bar{c}\bar{c}bb$, and $bb\bar{b}\bar{b}$ systems.